OpenStudy (anonymous):

find the dimensions of the rectangle. The perimeter is 30 meters and the length is 3 meters greater than the width.

OpenStudy (anonymous):

Let x be the width, then the length is x +3

OpenStudy (anonymous):

perimeter= 2 ( x + x +3)= 4 x +6=30 4x = 24 x=6=width length= 9